OpenStudy (anonymous):

Factor the GCF. 6m3 + 14m2 - 18m

OpenStudy (jdoe0001):

simplify each coefficient first say \(\bf 6 = 3\cdot 2\qquad \qquad 14=2\cdot 7\qquad \qquad 18=3\cdot 3\cdot 2\\ \quad \\ 6m^3 + 14m^2 - 18m\implies 3\cdot 2m^3+2\cdot 7m^2-3\cdot 3\cdot 2m\) can you see the common factor now?

OpenStudy (anonymous):

Oh, i think so! so the answer would be m(6m2 + 14m - 18)

OpenStudy (jdoe0001):

well, you have the variable correct...how about the coefficients? do they have a common factor too ?

OpenStudy (anonymous):

Wait ugh, now im getting 2m(3m2 + 7m - 9) am i even doing this right?

OpenStudy (jdoe0001):

\(\bf 3\cdot \color{blue}{2}m^3+\color{blue}{2}\cdot 7m^2-3\cdot 3\cdot \color{blue}{2}m \implies 2m(3m^2 + 7m - 9)\)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Oh i see, thank you very much!

OpenStudy (jdoe0001):

yw